# Mixture of Mixups for Multi-label Classification

### v1

@@ Abstract
Multi-label imbalanced classification poses a significant challenge in machine learning, particularly evident in bioacoustics where animal sounds often co-occur, and certain sounds are much less frequent than others. This paper focuses on the specific case of classifying anuran species sounds using the dataset AnuraSet, that contains both class imbalance and multi-label examples. To address these challenges, we introduce Mixture of Mixups (Mix2), a framework that leverages mixing regularization methods Mixup, Manifold Mixup, and MultiMix. Experimental results show that these methods, individually, may lead to suboptimal results; however, when applied randomly, with one selected at each training iteration, they prove effective in addressing the mentioned challenges, particularly for rare classes with few occurrences. Further analysis reveals that Mix2 is also proficient in classifying sounds across various levels of class co-occurrences.  [^1] Our code is released at: https:/github.com/ilyassmoummad/Mix2
@@ II-D Novel contribution: Mixture of Mixups (Mix2)
We now present Mix2, a probabilistic mixture of the three aforementioned methods. At each SGD training iteration, Mix2 employs random selection from three mixing regularization options: Mixup, Manifold Mixup, or Multimix, with probabilities of 25%, 50%, and 25%, respectively. This dynamic approach ensures that the model learns from a diverse range of augmented samples, promoting robustness and generalization.
@@ III-A Dataset
Given the absence of a dedicated validation set, designing one from the training set is challenging due to the overlap between examples. The split should be conducted at the audio recording level, from which examples were extracted using a sliding window. Additionally, due to the long-tail distribution, splitting classes with few occurrences into distinct subsets is difficult; this problem is already evident between the train and test sets. Consequently, creating a proper validation set poses a challenge. In line with standard machine learning practices, we conduct model validation on the test set.
@@ III-F Training
For all our experiments, we train MobileNetV3-Large [27] from scratch using AdamW optimizer with a batch size of 128, a learning rate of $10^{-2}$, and a weight decay of $10^{-6}$ for 100 epochs. We train all our systems for 100 epochs, with the exception of the first row (10 epochs) in Table I for a fair comparison with the baseline systems of AnuraSet, which are trained for 10 epochs.
@@ IV Results and Discussions
In Figure 3, we study the macro $F$-score score of examples with different polyphony levels ranging from 1 to 9 for Mix2, Mixup, and when no mixing is applied. We observe a similar trend where Mixup performs worse than when it is not applied, and Mix2 significantly boosts the performance. Error bars are shown at each point, with the standard deviation being very small across runs.

### v2

@@ Abstract
Multi-label imbalanced classification poses a significant challenge in machine learning, particularly evident in bioacoustics where animal sounds often co-occur, and certain sounds are much less frequent than others. This paper focuses on the specific case of classifying anuran species sounds using the dataset AnuraSet, that contains both class imbalance and multi-label examples. To address these challenges, we introduce Mixture of Mixups (Mix2), a framework that leverages mixing regularization methods Mixup, Manifold Mixup, and MultiMix. Experimental results show that these methods, individually, may lead to suboptimal results; however, when applied randomly, with one selected at each training iteration, they prove effective in addressing the mentioned challenges, particularly for rare classes with few occurrences. Further analysis reveals that the model trained using Mix2 is also proficient in classifying sounds across various levels of class co-occurrences.  [^1] Our code is released at: https:/github.com/ilyassmoummad/Mix2
@@ II-D Novel contribution: Mixture of Mixups (Mix2)
We now present Mix2, a probabilistic mixture of the three aforementioned methods. At each SGD training iteration, Mix2 employs random selection from three mixing regularization options: Mixup, Manifold Mixup, or Multimix. This dynamic approach ensures that the model learns from a diverse range of augmented samples, promoting robustness and generalization.
@@ III-A Dataset
Given the absence of a dedicated validation set, designing one from the training set is challenging due to the overlap between examples. The split should be conducted at the audio recording level, from which examples were extracted using a sliding window. Additionally, due to the long-tail distribution, splitting classes with few occurrences into distinct subsets is difficult; this problem is already evident between the train and test sets. Consequently, creating a proper validation set poses a challenge. In line with standard machine learning benchmarks, we conduct model validation on the test set.
@@ III-F Training
For all our experiments, we train MobileNetV3-Large [25] from scratch using AdamW optimizer with a batch size of 128, a learning rate of $10^{-2}$, and a weight decay of $10^{-6}$ for 100 epochs. We train all our systems for 100 epochs, with the exception of the first row (10 epochs) in Table I for a fair comparison with the baseline systems of AnuraSet, which are trained for 10 epochs. When we train the model using Mix2, we select Mixup, Manifold Mixup, or Multimix, with probabilities of 25%, 50%, and 25%, respectively. Manifold Mixup is selected more frequently as it was the best-performing single regularization method.
@@ IV Results and Discussions
In addition, we experimented with weighted sampling by oversampling rare combinations of classes. However, this approach yielded worse results than the baseline and was therefore excluded from the study.
In Figure 3, we study the macro $F$-score score of examples with different polyphony levels ranging from 1 to 9 for Mix2, Mixup, and when no mixing is applied. We observe a similar trend where Mixup performs worse than when it is not applied, and Mix2 noticeably improves the performance, a finding worthy of future investigation. Error bars are shown at each point, with the standard deviation being very small across runs.